# A Survey on Large Language Models for Critical Societal Domains: Finance, Healthcare, and Law

### v1

@@ 6.1.3 Domain-specific Ethics
On top of general ethical principles and considerations, in the specific context of different domains, the definitions of ethics display their distinct focus and subtle differences. Here, we introduce domain-specific investigations of ethics in finance, health, and law, respectively.
@@ 6.2 Future Prospects
For future work addressing ethical concerns in LLMs, a multifaceted strategy is imperative. (1) Dataset censorship: Meticulous dataset censorship is vital, involving a thorough examination and elimination of improper content (e.g., biased or incorrect information) from the training data. This step ensures that the model is shielded from potentially harmful information, minimizing the risk of encoding unwanted patterns. (2) Human and domain knowledge: The integration of humans in the AI loop is essential. Human reviewers contribute nuanced perspectives, provide domain knowledge, identify ethical issues, and guide the model’s learning process by refining its responses. Human-in-the-loop systems allow for ongoing monitoring and adjustments to address emerging ethical problems. (3) Theoretical bounds: Establishing theoretical bounds on the model’s behavior is important. The development of clear theoretical frameworks and ethical guidelines helps delineate the limits of the model’s decision-making, preventing it from generating potentially harmful or biased outputs. Through the implementation of these measures, we can elevate the ethical standards of LLMs, fostering responsible AI development. (4) Explanation and causality: It is essential to delve into the underlying causes and mechanisms behind the generation of LLM outputs. Understanding the root causes of ethical issues is beneficial for developing effective mitigation strategies.
@@ 7 Conclusion
Author Contributions
All authors contribute to paper writing.
Zhiyu Zoey Chen and Armineh Nourbakhsh lead the finance section.
Xinlu Zhang and An Yan lead the healthcare section.
Nan Hao leads the law section.
Jing Ma leads the ethics section.

### v2

@@ 6.1.3 Domain-specific Ethics
On top of general ethical principles and considerations, in the specific context of different domains, the definitions of ethics display their distinct focus and subtle differences. Here, we introduce domain-specific investigations of ethics in FHL domains, respectively.
@@ 6.2 Future Prospects
6.2.1 Methodological Directions
Looking ahead, the future prospects for LLM in FHL domains are at an exciting moment. Here, we outline several promising methodological directions that can be applied to address various ethical concerns:
  • Dataset censorship. Meticulous dataset censorship is vital, involving a thorough examination and elimination of improper content from the training data. This step ensures that the model is shielded from potentially harmful information, minimizing the risk of encoding unwanted patterns. For example, removing biased, private, or incorrect information in FHL scenarios can promote fairness, privacy, and reduction of hallucinations. Considering the specialized knowledge required in FHL domains, developing high-quality data censorship mechanisms presents significant challenges.
  • Human and domain knowledge for ethics. The integration of humans in the AI loop is essential. Human reviewers contribute nuanced perspectives, provide domain knowledge, identify ethical issues, and guide the model’s learning process by refining its responses. Human-in-the-loop systems allow for ongoing monitoring and adjustments to address emerging ethical problems. For example, as discussed in Section 5, involving legal advice from human expert would not only improve legal reasoning but also address complicated concerns in ethical and moral dilemmas.
  • Theoretical bounds. Establishing theoretical bounds on the model’s behavior is important. The development of clear theoretical frameworks and ethical guidelines helps delineate the limits of the model’s decision-making, preventing it from generating potentially harmful or biased outputs. Through the implementation of these measures, we can elevate the ethical standards of LLMs, fostering responsible AI development. For example, for the robustness issue we discussed in Section 6.1.2, certifiable approaches for LLMs are crucial to theoretically guarantee model safety against a range of adversarial attacks.
  • Causality-involved analysis. It is essential to delve into the underlying causes and mechanisms behind the generation of LLM outputs. For example, concerning the explainability and fairness issues discussed in Section 6.1.2, incorporating causality can help elucidate the causes behind model behavior and eliminate potential biases against underrepresented groups. By understanding and addressing the causal relationships within the data and the model, we can develop effective strategies to improve both the transparency and equity of the LLM outputs.
6.2.2 Further Ethical Concerns in FHL
Here, we highlight several urgent and critical concerns for LLM + FHL areas, which hold substantial potential for future development.
Safety and privacy. Employing LLMs in FHL sectors introduces substantial safety and privacy issues, given the sensitive nature of the information handled in these areas. Existing studies (ThankGod Chinonso, 2023) have reported LLM issues in privacy and safety, coming from both model-inherent vulnerabilities (e.g., data extraction, data poisoning) and other vulnerabilities (e.g., prompt injection), as we also introduced in Section 6.1. Traditional privacy and safety strategies often fail on LLMs due to the large scale of LLM parameters and high computation cost. Considering this, we list the following lines for future research: (1) Clear guidelines and user education: Provide clear, understandable guidelines that inform FHL users about the risks associated with the provision of personal information. Regularly remind and educate users on the importance of being cautious with the amount of personal information they share. Emphasize the use of minimal and relevant data inputs when interacting with LLMs to reduce privacy risks. (2) Vigilant third-party management: Implement a rigorous screening process for all third-party partners who may have access to the data used by LLMs. Establish strict data handling and confidentiality agreements with third-party partners to prevent data leakage. These agreements should enforce adherence to privacy standards and include penalties for non-compliance. (3) Robust defense strategies for LLMs in FHL: Develop a FHL domain-specific security framework for LLMs to protect against unauthorized access and data breaches. Regularly evaluate and update defense strategies to address emerging security threats, and establish a robust incident response plan to handle potential security breach.
Explainability. We have discussed the general considerations and methods for explanability in Section 6.1.2. In many settings, LLM applications in FHL can inherit or extend general LLM explanation methods. On top of that, in special FHL context that general explanation methods cannot tackle, some promising directions may include: (1) Domain-specific customization: Tailor explanations to the specific needs and comprehension levels of different stakeholders in FHL sectors, recognizing that financial experts, healthcare providers, and legal professionals may require different types of information to effectively trust and use AI outputs. (2) Feedback loops for continuous improvement: As FHL domains often involve long-term or multi-stage processes, we need to establish mechanisms for users to provide feedback on explanations, which can be used to continually refine explanation methodologies.
Potential exacerbation of existing inequalities. As the bias issue we discussed in Section 6.1, LLM applications in FHL areas may suffer from exacerbation of inequalities. We suggest the following directions for future studies: (1) Definition of FHL domain-specific bias and fairness notions: On top of general bias notions such as group fairness and individual fairness, it is crucial to first establish clear definitions of bias and fairness specifically in the FHL domains. These definitions should reflect the unique characteristics and requirements of each sector that have not been covered in existing general bias problems. (2) Generative FHL data to eliminate bias: One promising direction to combat bias is the generation of synthetic data that reflects the diversity and complexity of real-world scenarios in each FHL domain. By enriching training datasets with generative techniques, researchers can better model minority cases and underrepresented groups. (3) Legacy support and impact assessments: Integrate bias and fairness impact assessments as part of the compliance process for FHL-focused LLMs. Such assessments should be regularly updated and publicly accessible to maintain trust and accountability.
Job displacement. Widespread adoption of LLMs could lead to significant workforce changes and potential job losses. But it is important to consider that LLMs bring efficiency improvement for routine and entry-level tasks; meanwhile, they also present opportunities for workforce enhancement and the creation of new job roles. For instance, while routine tasks may be automated, there will be an increased demand for skilled professionals to manage, interpret, and oversee these technologies. This shift could lead to the emergence of new specialties and roles that are currently unforeseen, contributing positively to job creation in the economy. To address the potential for job changes, we suggest several potential proactive actions: (1) Upskilling and reskilling programs: Implement training programs to help current employees adapt to new roles that require managing or working alongside LLM technologies. (2) Policy development: Encourage the creation of policies that support workforce transitions. This includes social safety nets and incentives for businesses to retrain rather than replace their workforce. Foster collaboration between policymakers, educational institutions, industry leaders, and workers’ representatives to ensure a holistic approach to the challenges and opportunities posed by LLMs.
Accountability and liability. We suggest the following points to be considered for all stakeholders: (1) Clear attribution of responsibility: Define clear guidelines for attributing responsibility among all stakeholders involved in the development and deployment of LLMs, including developers, deployers, and end-users to help establish a chain of accountability. (2) Scenario-based guidance: Explore various scenarios where LLM outputs could lead to errors or harm and propose specific guidance on how accountability and liability should be managed in each case. This approach can help clearly prepare stakeholders for potential issues and accountability. (3) Stakeholder collaboration: Encourage collaboration among legal experts, ethicists, technologists, and policymakers to develop a comprehensive framework that addresses accountability and liability. This multidisciplinary approach is essential for creating robust solutions that are practical and enforceable.
Widening technology gaps. Smaller organizations and non-profits may face barriers to accessing advanced LLMs, potentially exacerbating disparities in service quality. Potential directions to address this issue may include: (1) Open-source initiatives: Promote the development and use of open-source LLM tools that could reduce costs and increase accessibility for all organizations. (2) Public-private partnerships: Facilitate partnerships between government bodies, academic institutions, and private sector leaders to provide the necessary resources and training to under-resourced groups. (3) Subsidized pricing models: Encourage commercial LLM developers to offer tiered pricing or subsidies for non-profits. (4) Policy recommendations: Propose specific policy interventions that could support wider access to these technologies. This might include government-funded programs to aid smaller entities in adopting LLMs or incentives to support non-profit and smaller organizations. (5) Awareness and training programs: Suggest the implementation of targeted programs to raise awareness and improve the AI literacy of smaller organizations and non-profits, enabling them to effectively use and benefit from LLM technologies.
Risk assessment of unintended consequences. LLM adoption could have unforeseen effects on FHL applications. We suggest the following strategies to handle unintended consequences: (1) Risk assessment and management: Develop systematic models for risk assessment that anticipate potential negative impacts of LLMs on FHL tasks. This includes predicting volatility, assessing the risk of bias, and identifying points of failure. (2) Continuous monitoring: Implement continuous monitoring systems that can dynamically assess and manage the risks associated with operational LLMs. These systems would use real-time data to update risk profiles and suggest necessary adjustments to deployment strategies. (3) Consequence explanability: Develop and enforce standards for explainability that require LLMs to provide understandable and detailed justifications for their outputs and potential outcomes at each step, so that developers can ensure that systems are easier to audit and improve over time.
Public trust and acceptance. Building public trust in AI systems is essential for their successful deployment. Potential future strategies to build and maintain public trust in LLMs may include: (1) Transparency initiatives: Make the algorithms’ decision-making processes more accessible and understandable to the general public, as well as disclose any limitations and uncertainties associated with AI predictions. (2) Educational outreach: Suggest strategies for educational outreach to improve public understanding of AI. This could include partnerships with educational institutions to integrate AI literacy into curriculums or public workshops and seminars that explain how LLMs are used in specific domains. (3) Stakeholder involvement: Recommend involving a broad range of stakeholders in the development and implementation of LLM policies. This should include policymakers, community leaders, consumer advocates, LLM developers, and other representatives from the society to ensure that multiple perspectives are considered and that AI deployments are aligned with public values and needs.